# On the combinatorics of tableaux — Classification of lattices underlying Schensted correspondences

### v1

@@ 5. Local characterization
Henceforth we will characterize admissible Fomin lattices under a certain set of assumptions. As we proceed through the analysis, additional, stricter assumptions will be made. This may seem an excessively narrow approach, but all known Fomin lattices with $\hat{0}$ other than $\mathbb{YF}_{r}$ satisfy all of these assumptions.
@@
Henceforth, we will restrict our attention to distributive Fomin lattices with $\hat{0}$. Specifically, we assume $L$ is a Fomin lattice with $\hat{0}$ with differential degree $r$, values $V$, and weighting $w$, and that $L$ is distributive. (We do not, at this point, assume that $w$ is admissible.)
@@
[Stan2012a]*Prop. 3.4.3[Birk1967a]*§III.3, Th. 3 Since we have assumed that $\hat{0}$ exists, $L$ is finitary, that is, all if its principal ideals are finite. $L$ is isomorphic to the set of finite order ideals of the poset $J=J(L)$ of its join-irreducible elements, which we denote $L={\operatorname{Ideal}_{\textup{f}}}(J)$. Given a covering $x\lessdot y$ in $L$, $x\setminus y$ (considered as the set difference of two ideals of $J$) contains exactly one element of $J$, and the projectivity requirement (condition (A)(4)) on $w$ reduces to: There is a weighting $w$ from $J$ to $V$, and for every $x\lessdot y$ in $L$, $w(x\lessdot y)=w(y\setminus x)$.
@@
Because the image of $w$ in $L$ is the same as the image of $w$ on $J$, we abuse language by conflating the two meanings of $w$ when we are dealing them as entire functions. In particular, each of them has all of its values positive iff the other does, and so “$w$ is admissible” is used to mean both.
@@
The elements of $J$ are called points or boxes.
@@
Regarding (1): If there was some $a\leq b$ in $J$ for which $[a,b]$ is infinite, then the set of all principal ideals of the elements in $[a,b]$, $\{[x]\,|\,a\leq x\leq b\}$ would be an infinite subset of $L$ with all of its elements $\geq[a]$ and $\leq[b]$. That contradicts that $L$ is locally finite.
Regarding (2): Similarly, if there was an $a\in J$ whose principal ideal $[a]$ in $J$ was infinite, then $\{[x]\,|\,x\leq b\}$ would be an infinite subset of $L$ with all of its elements $\leq[a]$. That contradicts that, given our assumptions, th. 5.2 shows that $L$ is finitary. ∎
@@
The lattice of diagrams is isomorphic to the lattice of finite ideals of points under union and intersection, the size of a diagram is the cardinality of its corresponding ideal (the number of points it contains), and the number of tableaux with a particular shape is the number of linear orderings of its points (considered as a subposet of $J$). The number of colorings $c(x)$ of a diagram $x$ is the product of the weights $w$ of its points: $c(x)=\prod_{a\in x}w(x)$.
@@
Given any diagram $x$, its insertion points are the points that are the minimal elements of $J\setminus x$, those which can be added to $x$ to make a diagram (whose size is 1 larger than the size of $x$). The deletion points of $x$ are the points that are the maximal elements of $x$, those which can be deleted from $x$ to make a diagram (whose size is 1 smaller than the size of $x$). For any diagram $x$, we define $D_{x}$ to be the set of deletion points of $x$, and $I_{x}$ to be the set of insertion points of $x$.
@@
$\sum_{a\in I_{x}}w(a)+r=\sum_{b\in D_{x}}w(b)\textup{\qquad\qquad for all $x\in{\operatorname{Ideal}_{\textup{f}}}(J)$}.$ (10)
@@
$r=\sum_{z\,|\,z\textup{ minimal in }J}w(z)$ (11)
In many cases $J$ has a unique minimal element, which we denote $\hat{0}_{J}$, and in that case, (11) reduces $w(\hat{0}_{J})=r$.
Given an ideal $x$ and a point $p$ that is an insertion point of $x$, that is, minimal in $J\setminus x$, we can construct the ideal ${x^{\prime}}=x\sqcup\{p\}$. Thus $J\setminus x=(J\setminus x^{\prime})\sqcup\{p\}$ and $p$ is a deletion point of $x^{\prime}$.
@@
$\sum_{a\in D_{x^{\prime}}\setminus D_{x}}w(a)-\sum_{b\in D_{x}\setminus D_{x^{\prime}}}w(b)=\sum_{a\in I_{x^{\prime}}\setminus I_{x}}w(a)-\sum_{b\in I_{x}\setminus I_{x^{\prime}}}w(b)\textup{\qquad\qquad for all $x\in{\operatorname{Ideal}_{\textup{f}}}(J)$, $p\in I_{x}$}.$ (12)
@@
    $I_{x^{\prime}}\setminus I_{x}$: The set of minimal elements of $J\setminus{x^{\prime}}$ that are not minimal elements of $J\setminus x$,
@@
    $I_{x}\setminus I_{x^{\prime}}$: The set of minimal elements of $J\setminus x$ that are not minimal elements of $J\setminus{x^{\prime}}$, which is $\{p\}$.
@@
Condition (9) (for every $x\in{\operatorname{Ideal}_{\textup{f}}}(J)$) is equivalent to the combination of (11) and (12) (for every $x\in{\operatorname{Ideal}_{\textup{f}}}(J)$ and $p\in I_{x}$).
@@
Regarding $\Leftarrow$: Since every $y\in L$ is the top of a finite chain with $\hat{0}$ as bottom, we can can prove this by induction on increasing length of the chain. The base case is a chain of length zero, which implies $y=\hat{0}$. In that case, (9) with $x=y=\hat{0}$ is equivalent to (11).
If the chain has length >0>0, let $y^{\prime}$ be the element below $y$ in the chain. By induction, (9) is true for $x=y^{\prime}$. Now consider (12) with $x=y^{\prime}$ and $p$ the unique element of $y\setminus y^{\prime}$. We can add these two and by the above derivation obtain (9) for $x=y$. ∎
@@
Given a point $p\in J$, define $C^{-}_{p}$, or simply $C^{-}$ when $p$ is understood, ($C^{-}$ for “downward covers”) as the set of all elements of $J$ which are covered by $p$. Similarly, define $C^{+}_{p}$, or simply $C^{+}$, ($C^{+}$ for “upward covers”) as the set of all elements of $J$ which cover $p$. Define $N_{p}$, or simply $N$, (NN for “neighborhood”) to be $\{x\in J\,|\,(\exists a\in C^{-},b\in C^{+})\,a\leq x\leq b\}$, the convex subset of $J$ generated by $C^{-}$ and $C^{+}$. Clearly $C^{+}\subset N$ and $C^{-}\subset N$. Define $S_{p}$, or simply $S$, ($S$ for “siblings”) to be $N_{p}\setminus C^{-}\setminus C^{+}\setminus\{p\}$, $N_{p}$ with $^{-}, C+C^{+}$, and $p$ deleted.
@@
By construction, every element of $C^{+}$ is >> every element of $C^{-}$, but no element of $C^{+}$ covers any element of $C^{-}$. Every element of $S$ is incomparable to $p$, >> some element of $C^{-}$, and << some element of $C^{+}$. If $J$ is graded, every element of $S$ has the same rank as $p$ and $S$ is an antichain. However, in general, $S$ need not be an antichain.
@@
Given an ideal $x$ and a point $p$ that is an insertion point of $x$, $C^{-}\subset x$ and $C^{+}\subset J\setminus x^{\prime}$.
@@
Given an ideal $x$ and a point $p$ that is an insertion point of $x$, define $S^{-}_{xp}=S^{-}=S\cap x$ and $S^{+}_{xp}=S^{+}=S\cap(J\setminus x^{\prime})$.
@@
Since $p\not\in S$, by construction, $S$ is the disjoint union of $S^{-}$ and $S^{+}$, and if $s^{-}\in S^{-}$ and $s^{+}\in S^{+}$ are comparable, then $s^{-}\leq s^{+}$. So $S^{-}$ is an ideal of $S$ and $S^{+}$ is a filter of $S$.
The generic relationships of elements of these sets are shown in fig. 3.
@@
    $D_{x}\setminus D_{x^{\prime}}$ is the set of maximal elements of $x$ which are covered by $p$. This is the same as the elements of $C^{-}$ which are not $\leq$ (or equivalently, not covered by) any element of $S^{-}$. We denote this property by $\not\leq S^{-}$.
@@
    $I_{x^{\prime}}\setminus I_{x}$ is the set of minimal elements of $J\setminus x^{\prime}$ which cover $p$. This is the same as the elements of $C^{+}$ which are not $\geq$ (or equivalently, do not cover) any element of $S^{+}$. We denote this property by $\not\geq S^{+}$.
@@
$w(p)-\sum_{a\in C^{-},\ a\not\leq S^{-}}w(a)$ $=\sum_{b\in C^{+},\ b\not\geq S^{+}}w(b)-w(p)$
$\sum_{a\in C^{-}_{p},\ x\not\leq S^{-}_{xp}}w(a)+\sum_{b\in C^{+}_{p},\ y\not\geq S^{+}_{xp}}w(b)$ $=2w(p)\textup{\qquad\qquad for all $x\in{\operatorname{Ideal}_{\textup{f}}}(J)$, $p\in I_{x}$}.$ (13)
@@
Looking at (13) as a set of linear equations in the values of $w$, we see that each equation is determined by choosing first an $x\in{\operatorname{Ideal}_{\textup{f}}}(J)$ and then a $p\in I_{x}$. We will now reorganize the statement of the set of linear equations (13), but indexed first by an arbitrary choice of $p\in J$ (which determines $C^{+}$ and $C^{-}$), and then by the values of $S^{+}$ and $S^{-}$ consistent with $p$.
@@
A particular equation in the set (13) is indexed first by an $x\in{\operatorname{Ideal}_{\textup{f}}}(J)$ and then a $\in I_{x}. pp$ determines $C^{+}_{p}$ and $C^{-}_{p}$, and $p$ and $x$ together determine $S^{-}_{xp}$ and $S^{+}_{xp}$. Necessarily, $S_{p}=S^{+}_{xp}\sqcup S^{+}_{xp}$, so setting $T^{-}=S^{-}_{xp}$ and $T^{+}=S^{+}_{xp}$ shows that the equation is in the set (14).
Conversely, a particular equation in the set (14) is indexed first by a $p\in J$ and then a partition of $S_{p}$ into an ideal $T^{-}$ and a filter $T^{+}$. Define $x=\{z\in J\,|\,z<p\}\sqcup T^{-}=(p)\sqcup T^{-}$, the principal ideal of $J$ generated by $p$, less $p$ itself, plus $T^{-}$. By construction, $x$ is an ideal of $J$, $p$ is in insertion point of $x$, $S^{-}_{xp}=T^{-}$ and $S^{+}_{xp}=T^{+}$. This shows that the equation is in the set (13). ∎
@@
Define $x\in C^{+}$ as an upward orphan (relative to $p$) if $x$ does not cover any element of $S$, which is equivalent to $x$ covering only one element $n$ in $N$ (namely $p$). Similarly, we define $x\in C^{-}$ as a downward orphan (relative to $p$) if $x$ is not covered by any element of $S$, which is equivalent to $x$ being covered by only one element $n$ in $N$ (namely $p$).
In (14), if we choose $T^{-}=\emptyset$ and $T^{+}=S_{p}=S$, it becomes
$\sum_{a\in C^{-}}w(a)+\sum_{b\in C^{+},\ b\not\geq S}w(b)$ $=2w(p)\textup{\qquad for all $p\in J$}$
By construction, $S$ contains every element covered by any element of $C^{+}$ except for $p$, so the second sum is over all upward orphans. Thus
@@
$\sum_{a\,|\,a\lessdot p}w(a)+\sum_{b\,|\,b\gtrdot p,\ b\textup{ upward orphan}}w(b)$ $=2w(p)\textup{\qquad for all $p\in J$}$ (15)
@@
$\sum_{a\,|\,a\lessdot p,\ a\textup{ downward orphan}}w(a)+\sum_{b\,|\,b\gtrdot p}w(b)$ $=2w(p)\textup{\qquad for all $p\in J$}$ (16)
@@
Given $p$, let us partition $S_{p}$ into three disjoint parts $S_{p}=L\sqcup\{m\}\sqcup U$ where $L$ is an ideal of $S$, $U$ is a filter of $S$, and $m$ is an element between $L$ and $U$. Setting $E^{-}=L$ and $E^{+}=\{m\}\sqcup U$ in (14) and then setting $F^{-}=L\sqcup\{m\}$ and $F^{+}=U$ in (14) gives
$\sum_{a\,|\,a\in C^{-},\ a\not\leq E^{-}}w(a)+\sum_{b\,|\,b\in C^{+},\ b\not\geq E^{+}}w(b)$ $=2w(p)\textup{\qquad and}$
$\sum_{a\,|\,a\in C^{-},\ a\not\leq F^{-}}w(a)+\sum_{b\,|\,b\in C^{+},\ b\not\geq F^{+}}w(b)$ $=2w(p)\textup{\qquad for all $p\in J$ and $S_{p}=L\sqcup\{m\}\sqcup U$}$
Since $E^{-}\subset F^{-}$, $x\not\leq F^{-}$ implies $x\not\leq E^{-}$. And since $F^{+}\subset E^{+}$, $x\not\geq E^{+}$ implies $x\not\geq F^{+}$. Thus subtracting these two equations gives
$\sum_{a\,|\,a\in C^{-},\ a\not\leq E^{-},\ \textup{not }a\not\leq F^{-}}w(a)$ $=\sum_{b\,|\,b\in C^{+},\ b\not\geq F^{+},\ \textup{not }b\not\geq E^{+}}w(b)\textup{\qquad for all $p\in J$ and $S_{p}=L\sqcup\{m\}\sqcup U$}$
@@
Given a $p\in J$, (14) holds for every partition of $S_{p}$ into an ideal $T^{-}$ and a filter $T^{+}$ iff the following conditions are true for $p$:
@@
    (17) holds for every partition of $S_{p}$ into three disjoint sets $S_{p}=L\sqcup\{m\}\sqcup U$, where L is an ideal of $S_{p}$ and $U$ is a filter of $S_{p}$.
@@
Regarding $\Rightarrow$: This has been proved by the above derivations of (15) and (17).
@@
$\sum_{a\,|\,a\in C^{-}_{p},\ a\not\leq L,\ a\leq m}w(a)$ $=\sum_{b\,|\,b\in C^{+}_{p},\ b\not\geq U,\ b\geq m}w(b)$ (19)
Subtracting the left side of (19) from the first term of (18) and adding the right side of (19) from the second term of (18) gives
$\sum_{a\in C^{-}_{p},\ a\not\leq L,\ a\not\leq m}w(a)+\sum_{b\in C^{+}_{p},\ b\not\geq U}w(b)$ $=2w(p)$
This equation is the same as (14) for the partition $S_{p}=T^{-}\sqcup T^{+}$:
$\sum_{a\in C^{-}_{p},\ a\not\leq T^{-}}w(a)+\sum_{b\in C^{+}_{p},\ b\not\geq T^{+}}w(b)$ $=2w(p)$
Thus by induction (14) holds for every $p$ and partition of $S_{p}$ into an ideal $T^{-}$ and a filter $T^{+}$. ∎
Unique-cover-modular

### v2

@@ 5. Local characterization
We start with proving various “local” facts about the poset $P$ of points (join-irreducible elements) of $L$. These culminate with a classification of the possible structures of the neighborhood of a point $p$, that is, the convex set generated by the points that cover of $p$ and the points that are covered by $p$.
Henceforth we will characterize positive Fomin lattices under certain sets of assumptions. As we proceed through the analysis, further assumptions will be added. All of these assumptions may seem an excessively narrow approach, but all known Fomin lattices with $\hat{0}$ other than $\mathbb{YF}_{r}$ satisfy all of these assumptions.
@@
Henceforth, we will restrict our attention to distributive Fomin lattices with $\hat{0}$. Specifically, we assume $L$ is a Fomin lattice with $\hat{0}$ with differential degree $r$, values $V$, and weighting $w$, and that $L$ is distributive. (We do not, at this point, assume that $w$ is positive.)
@@
[Birk1967a]*§III.3 Th. 3[Stan2012a]*Prop. 3.4.3 Since we have assumed that $\hat{0}$ exists, $L$ is finitary, that is, all if its principal ideals are finite. $L$ is isomorphic to the set of finite order ideals of the poset $P=J(L)$ of its join-irreducible elements, which we denote $L={\operatorname{Ideal}_{\textup{f}}}(P)$. Given a covering $x\lessdot y$ in $L$, $x\setminus y$ (considered as the set difference of two ideals of $P$) contains exactly one element of $P$, and the projectivity requirement (condition $(A)$) on $w$ can be simplified to: There is a weighting function $w$ from $P$ into $V$, and for every $x\lessdot y$ in $L$, $w(x\lessdot y)=w(y\setminus x)$.
@@
Because the image of $w$ in $L$ is the same as the image of $w$ on $P$, we abuse language by conflating the two meanings of $w$ when we are dealing them as entire functions. In particular, each of them has all of its values positive integers iff the other does, and so “$w$ is positive” is used to mean both.
@@
The elements of $P$ are called points or boxes.
@@
Regarding (1): If there was some $a\leq b$ in $P$ for which $[a,b]$ is infinite, then the set of all (closed) principal ideals of the elements in $[a,b]$, $\{[x]\,|\,a\leq x\leq b\}$, would be an infinite subset of $L$ with all of its elements $\geq[a]$ and $\leq[b]$. That contradicts that $L$ is locally finite.
Regarding (2): Similarly, if there was an $a\in P$ whose principal ideal $[a]$ in $P$ was infinite, then $\{[x]\,|\,x\leq b\}$ would be an infinite subset of $L$ with all of its elements $\leq[a]$. That contradicts that, given our assumptions, th. 5.2 shows that $L$ is finitary. ∎
@@
The lattice of diagrams is isomorphic to the lattice of finite ideals of points under union and intersection, the size of a diagram is the cardinality of its corresponding ideal (the number of points it contains), and the number of tableaux with a particular shape is the number of linear orderings of its points (a subposet of $P$). The number of colorings $c(x)$ of a diagram $x$ is the product of the weights $w$ of its points: $c(x)=\prod_{a\in x}w(a)$.
@@
Given any diagram $x$, its insertion points are the points that are the minimal elements of $P\setminus x$, those which can be added to $x$ to make a diagram (whose size is 1 larger than the size of $x$). The deletion points of $x$ are the points that are the maximal elements of $x$, those which can be deleted from $x$ to make a diagram (whose size is 1 smaller than the size of $x$). For any diagram $x$, we define $D_{x}$ to be the set of deletion points of $x$, and $I_{x}$ to be the set of insertion points of $x$.
@@
$\sum_{a\in D_{x}}w(a)+r=\sum_{b\in I_{x}}w(b)\textup{\qquad\qquad for all $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$}.$ (10)
@@
$r=\sum_{z\,|\,z\textup{ minimal in }P}w(z)$ (11)
In many cases $P$ has a unique minimal element, $\hat{0}_{P}$, and in that case, (11) reduces to $w(\hat{0}_{P})=r$.
Given an ideal $x$ and a point $p$ in $I_{x}$, that is, minimal in $P\setminus x$, we can construct the ideal ${x^{\prime}}=x\sqcup\{p\}$. Thus $P\setminus x=(P\setminus x^{\prime})\sqcup\{p\}$ and $p\in D(x^{\prime})$.
@@
$\sum_{a\in D_{x^{\prime}}\setminus D_{x}}w(a)-\sum_{b\in D_{x}\setminus D_{x^{\prime}}}w(b)=\sum_{a\in I_{x^{\prime}}\setminus I_{x}}w(a)-\sum_{b\in I_{x}\setminus I_{x^{\prime}}}w(b)\textup{\qquad\qquad for all $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$, $p\in I_{x}$}.$ (12)
@@
    $I_{x^{\prime}}\setminus I_{x}$: The set of minimal elements of $P\setminus{x^{\prime}}$ that are not minimal elements of $P\setminus x$,
@@
    $I_{x}\setminus I_{x^{\prime}}$: The set of minimal elements of $P\setminus x$ that are not minimal elements of $P\setminus{x^{\prime}}$, which is $\{p\}$.
@@
Condition (9) (for every $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$) is equivalent to the combination of (11) and (12) (for every $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$ and $p\in I_{x}$).
@@
Regarding $\Leftarrow$: For any $y\in{\operatorname{Ideal}_{\textup{f}}}(P)$, we want to prove (9) for its $x=y$. Since $y$ is finite, we can prove this by induction upward on the size of $y$. The base case is $y=\emptyset=\hat{0}$. In that case, (9) with its $x=\hat{0}$ reduces to (11).
If $y$ has positive size, let $y^{\prime}\lessdot y$ in ${\operatorname{Ideal}_{\textup{f}}}(P)$. By induction, (9) is true for its $x=y^{\prime}$. Now consider (12) with its $x=y^{\prime}$ and its $p$ the unique element of $y\setminus y^{\prime}$. We can add these two equations and by the above derivation obtain (9) for its $x=y$. ∎
@@
Given a point $p\in P$, define $C^{-}_{p}$ ($C^{-}$ for “downward covers”) as the set of all elements of $P$ which are covered by $p$. Similarly, define $C^{+}_{p}$ ($C^{+}$ for “upward covers”) as the set of all elements of $P$ which cover $p$. Define $N_{p}$ (NN for “neighborhood”) to be $\{x\in P\,|\,(\exists a\in C_{p}^{-},b\in C_{p}^{+})\,a\leq x\leq b\}$, the convex subset of $P$ generated by $C_{p}^{-}$ and $C_{p}^{+}$. Define $S_{p}$ ($S$ for “siblings”) to be $\{x\in P\,|\,x\neq p\textup{ and }(\exists a\in C_{p}^{-},b\in C_{p}^{+})\,a<x<b\}=N_{p}\setminus C_{p}^{-}\setminus C_{p}^{+}\setminus\{p\}$, $N_{p}$ with $C_{p}^{-}$, $C_{p}^{+}$, and $p$ deleted.
@@
By construction, every element of $C_{p}^{+}$ is >> every element of $C_{p}^{-}$, but no element of $C_{p}^{+}$ covers any element of $C_{p}^{-}$. All elements of $C_{p}^{+}$ are incomparable to each other. Dually for $C_{p}^{-}$. Every element of $S_{p}$ is incomparable to $p$. If $P$ is graded, every element of $S_{p}$ has the same grade as $p$ and $S_{p}$ is an antichain. However, in general, $S_{p}$ need not be an antichain.
@@
Given an ideal $x$ and a point $p$ that is an insertion point of $x$, $C_{p}^{-}\subset x$ and $C_{p}^{+}\subset P\setminus(x\sqcup\{p\})$.
@@
Given an ideal $x$ and a point $p$ in $I_{x}$, define $S^{-}_{px}=S_{p}\cap x$ and $S^{+}_{px}=S_{p}\cap(P\setminus(x\sqcup\{p\}))$.
@@
Since $p\not\in S$, by construction, $S_{p}=S_{px}^{-}\sqcup S_{px}^{+}$, and if $s^{-}\in S_{px}^{-}$ and $s^{+}\in S_{px}^{+}$ are comparable, then $s^{-}\leq s^{+}$. So $S_{px}^{-}$ is an ideal of $S_{p}$ and $S_{px}^{+}$ is a filter of $S_{p}$.
The generic relationships of the elements of these sets are shown in fig. 7.
@@
Unique-cover-modular
Assumption 5.13.
Remark 5.14.
The poset $t$-Plait in [Fom1994a]*Exam. 2.2.10 and Fig. 7 is locally finite and cover-modular but is not graded.131313“Graded” means that every saturated chain from $a$ to $b$ has the same length, but there need not be a $\hat{0}$. In general, the grade function of a graded poset ranges over the integers, not just the nonnegative integers.
Lemma 5.15.
Let $p\in P$. For any $a\in C_{p}^{+}$ and any $b\neq p$ that is covered by $a$, unique-cover-modularity shows that there is a $c\in C_{p}^{-}$ that is covered by $a$ and so $b\in S_{p}$. Dually, for any $a\in C_{p}^{-}$ and any $b\neq p$ that covers $a$, there is a $c\in C_{p}^{+}$ that covers $a$ and so $b\in S_{p}$. For any $a\in S_{p}$, unique-cover-modularity shows there is a unique element of $C_{p}^{+}$ which is >a>a. Dually, there is a unique element of $C_{p}^{-}$ which is <a<a.
Definition 5.16.
For any $p\in P$ and any $a\in S_{p}$, we define $a^{+}_{p}$ to be the unique element of $C_{p}^{+}$ which is >a>a. Dually, we define $a^{-}_{p}$ to be the unique element of $C_{p}^{-}$ which is <a<a.
@@
    $D_{x}\setminus D_{x^{\prime}}$ is the set of maximal elements of $x$ which are covered by $p$. This is the same as the elements of $C_{p}^{-}$ which are not $\leq$ (or equivalently, not covered by) any element of $S_{px}^{-}$. We denote this property by $\not\leq S_{px}^{-}$.
@@
    $I_{x^{\prime}}\setminus I_{x}$ is the set of minimal elements of $P\setminus x^{\prime}$ which cover $p$. This is the same as the elements of $C_{p}^{+}$ which are not $\geq$ (or equivalently, do not cover) any element of $S_{px}^{+}$. We denote this property by $\not\geq S_{px}^{+}$.
@@
$w(p)-\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq S_{px}^{-}}w(a)$ $=\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq S_{px}^{+}}w(b)-w(p)$ for all $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$, $p\in I_{x}$
$\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq S^{-}_{px}}w(a)+\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq S^{+}_{px}}w(b)$ $=2w(p)$ $\textup{for all $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$, $p\in I_{x}$}.$ (13)
Proof.
Regarding (1) and (4): These are straightforward.
Regarding (2) $\Rightarrow$: Let $a\in D_{x}\setminus D_{x^{\prime}}$. Then $a$ is maximal in $x$ but not maximal in $x^{\prime}$}. Thus $a\leq$ an element of $x^{\prime}$ which is not an element of $x$. The only element of $x^{\prime}$ that is not an element of $s$ is $p$, so $a<p$. If $a$ is not covered by $p$, then there is a $b$ with $a<b<p$. Thus $b\in x^{\prime}$ but since $a$ is maximal in $x$, $b\not\in x$, which requires that $b=p$, which is a contradiction. Thus $a$ is covered by $p$ and so $a\in C^{-}_{p}$.
To show that $a\not\leq S_{px}^{-}$, assume there is a $c$ in $S_{px}^{-}$ with $a\leq c$. Since $C^{-}_{p}$ is disjoint from $S^{-}_{ps}$, $a<c$. But $c\in x$, contradicting that $a$ is maximal in $x$. Thus $a\not\leq S_{px}^{-}$.
Regarding (2) $\Leftarrow$: Let $a\in C^{-}_{p}$ and $a\not\leq S^{-}_{px}$. Because $a<p$ and $p\in x^{\prime}$, $a$ is not maximal in $x^{\prime}$.
To show that $a$ is maximal in $x$, assume there is a $d$ in $x$ with $d>a$. We can assume that $d\gtrdot a$. Then we apply unique-cover-modularity to $p$ and $d$ to obtain a $e$ which covers $p$ and $d$. So $e\in C^{+}_{p}$, and with $a\in C^{-}_{p}$ shows $d\in S_{p}$. Since $d\in x$, $d\in S^{-}_{px}$. This contradicts that $a\not\leq S^{-}_{ps}$, so $a$ is maximal in $x$.
Regarding (3): This is proved dually to (2). ∎
@@
Looking at (13) as a set of linear equations in the values of $w$, we see that each equation is determined by choosing first an $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$ and then a $p\in I_{x}$. We will now reorganize the statement of the set of linear equations (13), but indexed first by an arbitrary choice of $p\in P$ (which determines $C_{p}^{+}$ and $C_{p}^{-}$), and then by the values of $S_{px}^{+}$ and $S_{px}^{-}$ consistent with $p$.
@@
A particular equation in the set (13) is indexed first by an $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$ and then a $\in I_{x}. pp$ determines $C_{p}^{+}$ and $C_{p}^{-}$, and $p$ and $x$ together determine $S^{-}_{px}$ and $S^{+}_{px}$. Necessarily, $S_{p}=S^{+}_{px}\sqcup S^{+}_{px}$, so setting $T^{-}=S^{-}_{px}$ and $T^{+}=S^{+}_{px}$ shows that the equation is in the set (14).
Conversely, a particular equation in the set (14) is indexed first by a $p\in P$ and then a partition of $S_{p}$ into an ideal $T^{-}$ and a filter $T^{+}$. Define $x=\{z\in P\,|\,z<p\}\sqcup T^{-}=(p)\sqcup T^{-}$, the principal ideal of $P$ generated by $p$, less $p$ itself, plus $T^{-}$. By construction, $x$ is an ideal of $P$, $p$ is in insertion point of $x$, $S^{-}_{px}=T^{-}$ and $S^{+}_{px}=T^{+}$. This shows that the equation is in the set (13). ∎
@@
Define $x\in C_{p}^{+}$ as an upward orphan (relative to $p$) if $x$ does not cover any element of $S_{p}$, which is equivalent to $x$ covering only one element of $N_{p}$ (namely $p$). Similarly, we define $x\in C_{p}^{-}$ as a downward orphan (relative to $p$) if $x$ is not covered by any element of $S_{p}$, which is equivalent to $x$ being covered by only one element of $N_{p}$ (namely $p$).
In (14), if we choose $T^{-}=\emptyset$ and $T^{+}=S_{p}$, it becomes
$\sum_{a\in C_{p}^{-}}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq S_{p}}w(b)$ $=2w(p)\textup{\qquad for all $p\in P$}$
By lem. 5.15, $S_{p}$ contains every element covered by any element of $C_{p}^{+}$ except for $p$, so the second sum is over all upward orphans. Thus
@@
$\sum_{a\,|\,a\lessdot p}w(a)+\sum_{b\,|\,b\gtrdot p,\ b\textup{ upward orphan}}w(b)$ $=2w(p)\textup{\qquad for all $p\in P$}$ (15)
@@
$\sum_{a\,|\,a\lessdot p,\ a\textup{ downward orphan}}w(a)+\sum_{b\,|\,b\gtrdot p}w(b)$ $=2w(p)\textup{\qquad for all $p\in P$}$ (16)
@@
Given $p\in P$, let us partition $S_{p}$ into three disjoint parts $S_{p}=L\sqcup\{m\}\sqcup U$ where $L$ is an ideal of $S$, $U$ is a filter of $S$, and thus $m$ is an element between $L$ and $U$. In (14), setting its $T^{-}=L$ and its $T^{+}=\{m\}\sqcup U$ gives
$\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq L}w(a)+\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq\{m\}\sqcup U}w(b)$ $=2w(p)$ (17)
and then setting its $T^{-}=L\sqcup\{m\}$ and its $T^{+}=U$ in (14) gives
$\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq L\sqcup\{m\}}w(a)+\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq U}w(b)$ $=2w(p)$ (18)
Since $x\not\leq L\sqcup\{m\}$ implies $x\not\leq L$ and $x\not\geq\{m\}\sqcup U$ implies $x\not\geq U$, subtracting (18) from (17) gives
$\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq L,\ \textup{not }a\not\leq L\sqcup\{m\}}w(a)$ $=\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq U,\ \textup{not }b\not\geq\{m\}\sqcup U}w(b)$
@@
Given a $p\in P$, equation (14) holds for every partition of $S_{p}$ into an ideal $T^{-}$ and a filter $T^{+}$ iff:
@@
    equation (19) holds for $p$ and for every partition of $S_{p}$ into three disjoint sets $S_{p}=L\sqcup\{m\}\sqcup U$, where L is an ideal of $S_{p}$ and $U$ is a filter of $S_{p}$.
@@
Regarding $\Rightarrow$: This has been proved by the above derivations of (15), (16), and (19).
@@
$\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq L,\ a\leq m}w(a)$ $=\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq U,\ b\geq m}w(b)$ (21)
Subtracting the left side of (21) from the first term of (20) and adding the right side of (21) to the second term of (20) gives
$\sum_{a\in C_{p}^{-},\ a\not\leq L,\ a\not\leq m}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq U}w(b)$ $=2w(p)$
This equation is the same as (14) for the partition $S_{p}=(L\sqcup\{m\})\sqcup U=T^{-}\sqcup T^{+}$.
Thus by induction (14) holds for $p$ and every partition of $S_{p}$ into an ideal $T^{-}$ and a filter $T^{+}$. ∎
We can replace (15) with (16):
Lemma 5.23.
Given a $p\in P$, equation (14) holds for every partition of $S_{p}$ into an ideal $T^{-}$ and a filter $T^{+}$ iff:
  1. (1)
    equation (16) holds for $p$ and
  2. (2)
    equation (19) holds for $p$ and for every partition of $S_{p}$ into three disjoint sets $S_{p}=L\sqcup\{m\}\sqcup U$, where L is an ideal of $S_{p}$ and $U$ is a filter of $S_{p}$.
Proof.
This is proved dually to the proof of lem. 5.22, using (16) in place of (15). ∎
Positive weighting